miguel finds the equation of a line that is perpendicular to y=ax+b

. If he multiplies the slopes of the two lines, what is the product?


1
Expert's answer
2021-06-03T12:20:56-0400

Slope of a line y=mx+c is m

Equation of a perpendicular line to the line ax+by+c =0

is bx-ay+k=0

Given the equation of a line is:-

y=ax+b

"\therefore" Slope of this line =a

Now,

y=ax +b

"\implies" ax-y+b=0

"\therefore" General equation of a perpendicular line to ax-y+b=0 is ,

-x-ay+k=0

"\implies" -ay=x-k

"\implies y=-\frac{1}{a}x+\frac{k}{a}"

"\therefore Slope of the perpendicular line =-\frac{1}{a}"

Hence, the product of the slope ="a \cdot( \frac{-1}{a})"

=-1